# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ V-C3 Proof of Lemma 2
where $S_{0}$ and $\hat{S}_{0}$ are the same as matrices $S$ and $\hat{S},$ respectively, but with diagonal elements set to 00. To bound $\|D\|_{2},$ Observe that
$D_{ii}$ $=S_{ii}-\hat{S}_{ii}=\big{(}1-\frac{1}{d}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}p_{ij}\big{)}-\big{(}1-\frac{1}{d}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}\hat{p}_{ij}\big{)}$
$=\frac{1}{kd}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}\sum_{m=1}^{k}(Z_{m,{ij}}-p_{ij}).$ (79)
Also, since $D$ is a diagonal matrix, we have $\|D\|_{2}=\max_{i}|D_{ii}|$. And by Eq. 79, for any fixed $i$, $kdD_{ii}$ is a sum of at most $kd_{\max}$ independent, zero-mean random variables and each random variable takes values in $(-1,1)$. Therefore, by applying Hoeffding’s inequality, we have
$\forall i\in[n],\ {\mathbb{P}}\big{(}kd|D_{ii}|>t\big{)}$ $\leq 2\exp\bigg(-\frac{2t^{2}}{4kd_{\max}}\bigg{missing}).$ (80)
@@
${\mathbb{P}}\left(\|D\|_{2}\geq 3\sqrt{\frac{d_{\max}\log n}{kd^{2}}}\right)$ $\leq\sum_{i=1}^{n}{\mathbb{P}}\Big{(}|D_{ii}|>3\sqrt{\frac{d_{\max}\log n}{kd^{2}}}\Big{)}\leq O(n^{-3}).$ (81)
Bounding $\|S_{0}-\hat{S}_{0}\|_{2}$: The bound follows by a direct application of Lemma 9 on $\|kd(S_{0}-\hat{S}_{0})\|_{2}$. Substituting $t=2\sqrt{kd_{\max}\log n}$ utilizing our assumption that $d_{\max}\geq\log n$, we obtain
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a matrix such that $A_{ij}\sim\textup{Bin}(p_{ij},k)$ for every $(i,j)\in{\mathcal{E}}$ with $i\neq j$ and 00 otherwise. Then, we have
@@
Proof: We begin by representing $A_{ij}$ as a sum of $k$ independent Bernoulli random variables for $(i,j)\in{\mathcal{E}}$ and $i\neq j$. Specifically, let $A_{ij}=\sum_{m=1}^{k}Z_{m,{ij}},$ where $Z_{m,{ij}}\sim{\mathsf{Bernoulli}}(p_{ij})$. Also, define matrices $U^{ij},$ where
@@
Observe that $A-{\mathbb{E}}[A]=\sum_{(i,j)\in{\mathcal{E}},i\neq j}\sum_{m=1}^{k}U_{m}^{ij}$ and $\|A-{\mathbb{E}}[A]\|_{2}=\lambda_{\max}(\sum_{(i,j)\in{\mathcal{E}},i\neq j}\sum_{m=1}^{k}\tilde{U}_{m}^{ij})$. Now, we will employ the matrix Bernstein inequality to analyze the sum of independent, zero-mean self-adjoint random matrices. Clearly $\|\tilde{U}_{m}^{ij}\|_{2}\leq 1$, therefore we have
${\mathbb{P}}\bigg{(}\lambda_{\max}\bigg{(}\sum_{(i,j)\in{\mathcal{E}},i\neq j}\sum_{m=1}^{k}\tilde{U}_{m}^{ij}\bigg{)}\geq t\bigg{)}\leq 2n\exp\left(\frac{-t^{2}/2}{\sigma^{2}+t/3}\right),$
@@
$\sigma^{2}=\bigg{\|}\sum_{(i,j)\in{\mathcal{E}}:i\neq j}\sum_{m=1}^{k}{\mathbb{E}}[(\tilde{U}_{m}^{ij})^{2}]\bigg{\|}_{2}=\bigg{\|}\sum_{(i,j)\in{\mathcal{E}},i\neq j}\sum_{m=1}^{k}p_{ij}(1-p_{ij})\left(\begin{matrix}e_{i}e_{i}^{\mathrm{T}}&0\\ 0&e_{j}e_{j}^{\mathrm{T}}\end{matrix}\right)\bigg{\|}_{2}\leq\frac{kd_{\max}}{4}.$
@@ V-D Proof of Theorem 3
${\mathbb{P}}(T\geq t)$ $\leq{\mathbb{P}}\bigg{(}T_{3}\geq\bigg{(}t-\frac{c_{\beta}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}-(c_{\alpha}+c_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k}\bigg{)}\bigg{)}+{\mathbb{P}}\bigg{(}T_{1}\geq c_{\alpha}\frac{n\|\pi\|^{2}_{\infty}}{k}\bigg{)}$
$\ \ \ \ \ \ \ \ \ \ \ +{\mathbb{P}}\bigg{(}T_{2}\geq\frac{c_{\beta}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}+c_{\gamma}\frac{n\|\pi\|^{2}_{\infty}}{k}\bigg{)}$
@@
Define a vector $z_{ij}\in\mathbb{R}^{k}$ such that $z_{ij}=\left[z_{1_{ij}},\dots,z_{k_{ij}}\right]^{\mathrm{T}}$ and a matrix $A\in\mathbb{R}^{k\times k}$ such that $A={\mathbf{1}}_{k}{\mathbf{1}}_{k}^{\mathrm{T}}-I$. Observe that
$k(k-1)(T_{3}^{ij}-{\mathbb{E}}[T_{3}^{ij}])$ $=\left(\pi_{i}+\pi_{j}\right)^{2}\left(z_{ij}^{\mathrm{T}}Az_{ij}-p_{ij}^{2}{\mathbf{1}}_{k}^{\mathrm{T}}A{\mathbf{1}}_{k}\right)-2(\pi_{i}+\pi_{j})\pi_{j}\left(z_{ij}-p_{ij}{\mathbf{1}}_{k}\right)^{\mathrm{T}}A{\mathbf{1}}_{k}.$
Utilizing the fact that
$(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})=z_{ij}^{\mathrm{T}}Az_{ij}-2p_{ij}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A{\mathbf{1}}_{k}-p_{ij}^{2}{\mathbf{1}}_{k}^{\mathrm{T}}A{\mathbf{1}}_{k},$
we obtain
$k(k-1)(T_{3}^{ij}-\mathbb{E}[T_{3}^{ij}])=(\pi_{i}+\pi_{j})^{2}$ $(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})$
Observe that under hypothesis $H_{0}$, the last term of the above equation reduces to zero. We will utilize Hanson-Wright inequality [82, Theorem 1.1] to bound the first term. Therefore, under hypothesis $H_{0}$ and for some constant $c$ we have
$\mathbb{P}\left(k(k-1)\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:i\neq j\end{subarray}}T_{3}^{ij}\geq{k}(k-1)t\right)$ $={\mathbb{P}}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:i\neq j\end{subarray}}(\pi_{i}+\pi_{j})^{2}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})\geq{k}(k-1)t\right)$
$\leq{\mathbb{P}}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:i\neq j\end{subarray}}4\|\pi\|_{\infty}^{2}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})\geq{k}(k-1)t\right)$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\exp\left(-c\min\bigg{\{}\frac{k^{2}(k-1)^{2}t^{2}}{\|\tilde{A}\|_{\mathrm{F}}^{2}\cdot 4\|\pi\|_{\infty}^{4}\cdot\left(\frac{1}{2}\right)^{4}},\frac{k(k-1)t}{\|\tilde{A}\|_{2}\cdot 2\|\pi\|_{\infty}^{2}\cdot\left(\frac{1}{2}\right)^{2}}\bigg{\}}\right)$
=$\stackrel{{\scriptstyle\zeta_{2}}}{{=}}\exp\bigg(-c\min\bigg{\{}\frac{4k(k-1)t^{2}}{|{\mathcal{E}}|\|\pi\|_{\infty}^{4}},\frac{2kt}{\|\pi\|_{\infty}^{2}}\bigg{\}}\bigg{missing}),$ (85)
where $\zeta_{1}$ follows by applying Hanson-Wright inequality to the concatenated matrix $\tilde{A}\in{\mathbb{R}}^{|{\mathcal{E}}|k\times|{\mathcal{E}}|k}$ with matrix $A$ along its digonal and utilizing the fact that the entrywise sub-Gaussian norm of $Z_{m,{ij}}-p_{ij}$ is upper bounded by $\frac{1}{2}$ and $\zeta_{2}$ follows because $\|\tilde{A}\|^{2}_{\mathrm{F}}=|{\mathcal{E}}|k(k-1)$ and $\|\tilde{A}\|_{2}=\|A\|_{2}=k-1$. Finally, utilizing $|{\mathcal{E}}|\leq nd_{\max}$, we obtain
$\mathbb{P}\left(T_{3}>t\right)\leq\exp\bigg(-c\min\bigg{\{}\frac{4k(k-1)t^{2}}{nd_{\max}\|\pi\|_{\infty}^{4}},\frac{2kt}{\|\pi\|_{\infty}^{2}}\bigg{\}}\bigg{missing}).$
To obtain tail bounds under hypothesis $H_{1}$, we again utilize Eq. 84 as
@@
$\mathbb{P}\bigg{(}T_{3}-{\mathbb{E}}[T_{3}]$ $\leq-t\bigg{)}\leq\exp\bigg(-c\min\bigg{\{}\frac{4k(k-1)t^{2}}{nd_{\max}\|\pi\|_{\infty}^{4}},\frac{2kt}{\|\pi\|_{\infty}^{2}}\bigg{\}}\bigg{missing})+\exp\left(\frac{-kt^{2}}{32\|\pi\|_{\infty}^{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}}\right).$
This completes the proof.
VI Proofs of Lower Bound and Stability
In this section, we will provide proofs of Theorem 4 and Proposition 9.
VI-A Proof of Theorem 4
We will apply the Ingster-Suslina method to establish a lower bound on the critical threshold [71]. Throughout the proof, we will assume that the induced graph is complete. Moreover, under the null hypothesis, we assume that the pairwise comparison matrix $P$ is fixed to be an all $1/2$ matrix, i.e.,

### v2

@@ V-C3 Proof of Lemma 2
where $S_{0}$ and $\hat{S}_{0}$ are the same as matrices $S$ and $\hat{S}$, respectively, but with diagonal elements set to 0. To bound $\|D\|_{2},$ Observe that
$D_{ii}$ $=S_{ii}-\hat{S}_{ii}$
$=\left(\!1-\frac{1}{d}\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}p_{ij}\!\right)\!-\!\left(\!1-\frac{1}{d}\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}\hat{p}_{ij}\!\right)$
$=\frac{1}{kd}\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}\sum_{m=1}^{k}(Z_{m,{ij}}-p_{ij})\,.$ (77)
Also, since $D$ is a diagonal matrix, we have $\|D\|_{2}=\max_{i}|D_{ii}|$. And by Eq. 77, for any fixed $i$, $kdD_{ii}$ is a sum of at most $kd_{\max}$ independent, zero-mean random variables and each random variable takes values in (1,1)(-1,1). Therefore, by applying Hoeffding’s inequality, we have
[…]
@@
${\mathbb{P}}\bigg(\|D\|_{2}\geq$ $\ 3\sqrt{\frac{d_{\max}\log n}{kd^{2}}}\bigg)$
$\leq\sum_{i=1}^{n}{\mathbb{P}}\Big(|D_{ii}|>3\sqrt{\frac{d_{\max}\log n}{kd^{2}}}\Big)\leq O(n^{-3}).$ (79)
Bounding $\|S_{0}-\hat{S}_{0}\|_{2}$: The bound follows by a direct application of Lemma 9 (see below) on $\|kd(S_{0}-\hat{S}_{0})\|_{2}$. Substituting $t=2\sqrt{kd_{\max}\log n}$ utilizing our assumption that $d_{\max}\geq\log n$, we obtain
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a matrix such that $A_{ij}\sim\textup{Bin}(p_{ij},k)$ for every $(i,j)\in{\mathcal{E}}$ with $i\neq j$ and 0 otherwise. Then, we have
@@
Proof.
We begin by representing $A_{ij}$ as a sum of $k$ independent Bernoulli random variables for $(i,j)\in{\mathcal{E}}$ and $i\neq j$. Specifically, let $A_{ij}=\sum_{m=1}^{k}Z_{m,{ij}},$ where $Z_{m,{ij}}\sim{\mathsf{Bernoulli}}(p_{ij})$. Also, define matrices $U^{ij},$ where
@@
Observe that $A-{\mathbb{E}}[A]=\sum_{(i,j)\in{\mathcal{E}},i\neq j}\sum_{m=1}^{k}U_{m}^{ij}$ and $\|A-{\mathbb{E}}[A]\|_{2}=\lambda_{\max}(\sum_{(i,j)\in{\mathcal{E}}:i\neq j}\sum_{m=1}^{k}\tilde{U}_{m}^{ij})$. Now, we will employ the matrix Bernstein inequality to analyze the sum of independent, zero-mean self-adjoint random matrices. Clearly $\|\tilde{U}_{m}^{ij}\|_{2}\leq 1$, therefore we have
${\mathbb{P}}\left(\lambda_{\max}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\sum_{m=1}^{k}\tilde{U}_{m}^{ij}\right)\geq t\right)\leq 2n\exp\left(\frac{-t^{2}/2}{\sigma^{2}+t/3}\right),$
@@
$\sigma^{2}$ $=\left\|\sum_{(i,j)\in{\mathcal{E}}:i\neq j}\sum_{m=1}^{k}{\mathbb{E}}[(\tilde{U}_{m}^{ij})^{2}]\right\|_{2}$
$=\left\|\sum_{(i,j)\in{\mathcal{E}}:i\neq j}\sum_{m=1}^{k}p_{ij}(1-p_{ij})\left(\begin{matrix}e_{i}e_{i}^{\mathrm{T}}&0\\ 0&e_{j}e_{j}^{\mathrm{T}}\end{matrix}\right)\right\|_{2}$
$\leq\frac{kd_{\max}}{4}.$
@@ V-D Proof of Theorem 3
First, we begin by deriving bounds on the probability of type I error. Similar to the proof of Theorem 2, we split the test statistic $T$ as $T=T_{1}+T_{2}+T_{3}$, where $T_{1},T_{2}$ and $T_{3}$ are defined in Eqs. 51, 52 and 53. To calculate the ${\mathbb{P}}(T\geq t)$ for $t\geq 0$, we use the following inequality
${\mathbb{P}}(T\geq t)\leq$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.\left.-(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k}\right)\right)$
$\ +{\mathbb{P}}\bigg(T_{1}\geq\tilde{c}_{\alpha}\frac{n\|\pi\|^{2}_{\infty}}{k}\bigg)$
$\ +{\mathbb{P}}\bigg(T_{2}\geq\frac{\tilde{c}_{\beta}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\tilde{c}_{\gamma}\frac{n\|\pi\|^{2}_{\infty}}{k}\bigg)$
@@
$k(k-1)\big(T_{3}^{ij}-\mathbb{E}[T_{3}^{ij}]\big)$
$=(\pi_{i}+\pi_{j})^{2}Z_{ij}(Z_{ij}-1)+\pi_{j}^{2}k(k-1)$
$\ \ \ \ \ \ \ -2(k-1)(\pi_{i}+\pi_{j})\pi_{j}Z_{ij}$
$\ \ \ \ \ \ \ -k(k-1)\left(\left(\pi_{i}+\pi_{j}\right)p_{ij}-\pi_{j}\right)^{2}$
$=\left(\pi_{i}+\pi_{j}\right)^{2}\left(\sum_{m=1}^{k}Z_{m,{ij}}\right)\left(\sum_{m=1}^{k}Z_{m,{ij}}-1\right)$
$\ \ \ \ \ \ \ -2(k-1)\left(\pi_{i}+\pi_{j}\right)\pi_{j}\left(\sum_{m=1}^{k}Z_{m,{ij}}\right)$
$\ \ \ \ \ \ \ -k(k-1)\left(\pi_{i}+\pi_{j}\right)^{2}p_{ij}^{2}+2k(k-1)\left(\pi_{i}+\pi_{j}\right)\pi_{j}p_{ij}$
$=\left(\pi_{i}+\pi_{j}\right)^{2}\bigg(\sum_{m=1}^{k}\sum_{\begin{subarray}{c}m^{\prime}\neq m,\\ m^{\prime}=1\end{subarray}}^{k}\left(Z_{m,{ij}}Z_{m^{\prime},{ij}}-p_{ij}^{2}\right)\bigg)$
$\ \ \ \ \ \ \ \ \ \ -2(k-1)\left(\pi_{i}+\pi_{j}\right)\pi_{j}\left(\sum_{m=1}^{k}\left(Z_{m,{ij}}-p_{ij}\right)\right).$
Define a vector $z_{ij}\in\mathbb{R}^{k}$ such that $z_{ij}=\left[Z_{1,{ij}},\dots,Z_{k,{ij}}\right]^{\mathrm{T}}$ and a matrix $A\in\mathbb{R}^{k\times k}$ such that $A={\mathbf{1}}_{k}{\mathbf{1}}_{k}^{\mathrm{T}}-I_{k}$. Observe that
$k(k-1)(T_{3}^{ij}-{\mathbb{E}}[T_{3}^{ij}])=\left(\pi_{i}+\pi_{j}\right)^{2}\left(z_{ij}^{\mathrm{T}}Az_{ij}-p_{ij}^{2}{\mathbf{1}}_{k}^{\mathrm{T}}A{\mathbf{1}}_{k}\right)$
$\ \ \ \ \ \ \ \ \ -2(\pi_{i}+\pi_{j})\pi_{j}\left(z_{ij}-p_{ij}{\mathbf{1}}_{k}\right)^{\mathrm{T}}A{\mathbf{1}}_{k}.$
Utilizing the fact that
$(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})$
$\ \ \ \ \ \ \ =z_{ij}^{\mathrm{T}}Az_{ij}-2p_{ij}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A{\mathbf{1}}_{k}-p_{ij}^{2}{\mathbf{1}}_{k}^{\mathrm{T}}A{\mathbf{1}}_{k},$
we obtain
$k(k-1)(T_{3}^{ij}-\mathbb{E}[T_{3}^{ij}])$
$=(\pi_{i}+\pi_{j})^{2}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})$
Observe that under hypothesis $H_{0}$, the last term of the above equation reduces to zero. We will utilize Hanson-Wright inequality [rudelson2013hanson, Theorem 1.1] to bound the first term. Therefore, under hypothesis $H_{0}$ and for some constant $c$ we have
$\mathbb{P}\left(k(k-1)\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}T_{3}^{ij}\geq{k}(k-1)t\right)$
$={\mathbb{P}}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\frac{(\pi_{i}+\pi_{j})^{2}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})}{k(k-1)}\geq t\right)$
$\leq{\mathbb{P}}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\frac{4\|\pi\|_{\infty}^{2}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})}{k(k-1)}\geq t\right)$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.\left.\frac{k(k-1)t}{\|\tilde{A}\|_{2}\cdot 2\|\pi\|_{\infty}^{2}\cdot\left(\frac{1}{2}\right)^{2}}\right\}\right)$
@@
where $\zeta_{1}$ follows by applying the Hanson-Wright inequality to the concatenated matrix $\tilde{A}\in{\mathbb{R}}^{|{\mathcal{E}}|k\times|{\mathcal{E}}|k}$ with matrix $A$ along its diagonal $|{\mathcal{E}}|$ times and utilizing the fact that the entrywise sub-Gaussian norm of $Z_{m,{ij}}-p_{ij}$ is upper bounded by $\frac{1}{2}$, and $\zeta_{2}$ follows because $\|\tilde{A}\|^{2}_{\mathrm{F}}=|{\mathcal{E}}|k(k-1)$ and $\|\tilde{A}\|_{2}=\|A\|_{2}=k-1$. Finally, utilizing $|{\mathcal{E}}|\leq nd_{\max}$, we obtain
[…]
To obtain tail bounds under hypothesis $H_{1}$, we again utilize Eq. 82 as follows:
$\mathbb{P}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}(T_{3}^{ij}-{\mathbb{E}}[T_{3}^{ij}])\leq-t\right)\leq$
$\mathbb{P}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}(\pi_{i}+\pi_{j})^{2}(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A(z_{ij}-p_{ij}{\mathbf{1}}_{k})\right$.
[…]
${\mathbb{P}}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}2(\pi_{i}+\pi_{j})((\pi_{i}+\pi_{j})p_{ij}-\pi_{j})(z_{ij}-p_{ij}{\mathbf{1}}_{k})^{\mathrm{T}}A{\mathbf{1}}_{k}\right$.
[…]
$\mathbb{P}\left(\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\sum_{m=1}^{k}(\pi_{i}+\pi_{j})((\pi_{i}+\pi_{j})p_{ij}-\pi_{j})(Z_{m,{ij}}-p_{ij})\right$.
[…]
$\ \ \ \ \ \ \ +\exp\left(\frac{-2kt^{2}/16}{4\|\pi\|_{\infty}^{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}}\right),$
@@
$\mathbb{P}\bigg(T_{3}-{\mathbb{E}}[T_{3}]$ $\leq-t\bigg)\leq$
[…]
$\exp\left(\frac{-kt^{2}}{32\|\pi\|_{\infty}^{2}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}}\right).$
This completes the proof. ∎
VI Proof of Lower Bound
In this section, we will provide the proof of Theorem 4.
Proof of Theorem 4.
We will apply the Ingster-Suslina method to establish a lower bound on the critical threshold [IngsterSuslina2003]. Throughout the proof, we will assume that the induced graph is complete. Moreover, under the null hypothesis, we assume that the pairwise comparison matrix $P$ is fixed to be an all $1/2$ matrix, i.e.,